# Emergent Semantics Beyond Token Embeddings: Transformer LMs with Frozen Visual Unicode Representations

### v1

@@ 5.3.   EMERGENT SEMANTICS VISUALIZATION OF INPUT TOKEN EMBEDDINGS
Figure 7: t-SNE visualization of input token embeddings. In our frozen, visual Unicode embeddings (bvv241-nemo), clusters are sharp and distinct, but reflect only surface or formal properties (such as token length), with no alignment to the semantic categories. This highlights that emergent meaning in our models must arise above the embedding layer.
This illustrates that any emergence of true semantic structure must occur entirely within the network’s compositional layers.
@@ 6   DISCUSSION
6.1.   RE-EVALUATING EMBEDDINGS: FROM MEANING CONTAINERS TO STRUCTURAL PRIMITIVES
@@
By contrast, our frozen visual embeddings resolve this conflict. They provide a stable, information-rich, and purely structural foundation. The model does not expend capacity learning what a token looks like; it only learns what to do with it. This frees the Transformer’s attention and MLP layers to focus exclusively on their core competency: composing these structural primitives into high-level meaning.
@@
6.2.   IMPLICATIONS FOR ARCHITECTURE AND EFFICIENCY
@@
  • Efficient MoE and Model Fusion: Mixture-of-Experts (MoE) models could share a common frozen embedding layer, reducing parameter count and simplifying routing logic.
@@
6.3.   LIMITATIONS AND FUTURE WORK

### v4

@@ 5.3 EMERGENT SEMANTICS VISUALIZATION OF INPUT TOKEN EMBEDDINGS AND DEEP TRANSFORMER LAYERS
Figure 7: t-SNE visualization of the frozen visual embeddings. The prominent clustering directly corresponds to token length. This phenomenon is a direct artifact of our embedding generation process: multi-character tokens, when rendered and resized to a fixed-resolution bitmap, produce images with higher average pixel density. As Principal Component Analysis (PCA) is designed to capture axes of maximum variance, it naturally identifies this structural feature as a dominant component. The resulting clusters are therefore based on a physical token property (length/density), not semantic content, reinforcing the conclusion that semantic understanding must emerge entirely within the subsequent Transformer layers.
Figure 8: t-SNE visualization of input token embeddings for baseline (’Model unfrozen’). In this trainable embeddings, semantic groups (e.g., numbers, professions, animals) tend to lie closer to one another and clustering.
Figure 9: t-SNE visualization of frozen input token embeddings for ’Model_UNI_GLIF’. In the this frozen embeddings, semantic groups (e.g., numbers, professions, animals) located randomly and without clustering.
The following part (Figures 8, 9, 10, 11) presents a method for analyzing internal representations of a language model through extraction and visualization of activations from various neural network layers. The study examined 70 words organized into seven semantic groups (royalty, animals, cities, actions, objects, feelings, professions). Individual words were fed as input to the model, and vector representations were extracted using forward hooks mechanism from key architectural components: input embeddings, attention projections (v_proj and o_proj), and transformer block outputs across all layers of the model.
Figure 10: t-SNE visualization of first two (layer 0 - layer 1) and the last two (layer 14 - layer 15) attention projections (v_proj and o_proj), and transformer block outputs for baseline (’Model unfrozen’). In this picture semantic groups tend to clustering.
The collected high-dimensional vectors were normalized using StandardScaler, followed by t-SNE algorithm application for dimensionality reduction to a two-dimensional space for visualization purposes. This approach enabled tracking the evolution of semantic representations as information propagates through the model layers. The analysis reveals how initially dispersed tokens gradually organize into semantically related clusters in deeper network layers, demonstrating the model’s capability for hierarchical extraction and structuring of semantic features. This approach provides a tool for interpreting the "black box" of transformer models and understanding the mechanisms of semantic representation formation during natural language processing.
Figure 11: t-SNE visualization of first two (layer 0 - layer 1) and the last two (layer 14 - layer 15) attention projections (v_proj and o_proj), and transformer block outputs for ’Model_UNI_GLIF’ with frozen input embedding layer. In this picture semantic groups (e.g., numbers, professions, animals) tend to make clusters after frozen embedding layer demonstrating emergence of semantic structure. This visualization provides direct evidence that semantic clustering, absent in the frozen input, is formed by the model’s compositional blocks.
5.4 ABLATION STUDY: THE ROLE OF EMBEDDING INITIALIZATION
This ablation study compares the performance of traditional trainable embeddings, frozen embeddings initialized with Unicode-based visual representations, and frozen embeddings initialized with random black-and-white bitmaps, which serve as a randomized baseline.
Figure 12: Learning curves for the baseline model ’Model unfrozen’ vs frozen $n$.Embedding layer model ’Model UNI_GLIF’ and randomly initialized frozen $n$.Embedding layer models
Figure 13: MMLU and ARC-e performance for the baseline model ’Model unfrozen’ vs frozen $n$.Embedding layer model ’Model UNI_GLIF’ and randomly initialized frozen $n$.Embedding layer models
Figure 14: ARC-c and C-SENSE performance for the baseline model ’Model unfrozen’ vs frozen $n$.Embedding layer model ’Model UNI_GLIF’ and randomly initialized frozen $n$.Embedding layer models
This ablation study reveals an insight: even when frozen embeddings are initialized from random noise, the model still converges (Figure 12, Figure 13, Figure 14). This finding suggests that the semantic structures observed in trainable embeddings are not a prerequisite for learning, but rather an emergent artifact. We hypothesize that this phenomenon stems from a trainable embedding layer being implicitly forced to solve two conflicting tasks: 1) Token Discrimination, the primary goal of assigning a unique vector to each token, and 2) Semantic Proximity, the secondary goal of clustering vectors for contextually similar tokens. These objectives are often at odds. In essence, during backpropagation, gradients from high-level contextual errors "imprint" semantic relationships onto the embedding vectors, forcing them to serve as a low-dimensional projection surface for knowledge learned by the entire network. Our proposed frozen-embedding approach resolves this conflict by decoupling the tasks. By providing a stable, unique identifier for each token, the fixed vectors (whether visually or randomly initialized) perfectly satisfy the discrimination objective. This liberates the Transformer’s compositional layers—the attention and feed-forward networks—to exclusively manage the learning of semantic relationships. We posit that this separation of concerns reduces optimization conflicts and allows the architecture to learn more efficiently, as evidenced by the superior performance on reasoning benchmarks.
@@ 6 DISCUSSION
6.1 RE-EVALUATING EMBEDDINGS: FROM MEANING CONTAINERS TO STRUCTURAL PRIMITIVES
@@
6.2 IMPLICATIONS FOR ARCHITECTURE AND EFFICIENCY
@@
  • Efficient MoE and Model Fusion: Mixture-of-Experts (MoE) models, which typically shard their MLP layers, could additionally benefit from a shared frozen embedding layer, reducing total parameter count and simplifying the routing logic by providing a common input space for all experts.
@@
6.3 LIMITATIONS AND FUTURE WORK